# Binomial Self-compensation for Motion Error in Dynamic 3D Scanning

### v1

@@ Abstract
Phase shifting profilometry (PSP) is favored in high-precision 3D scanning due to its high accuracy, robustness, and pixel-wise property. However, a fundamental assumption of PSP that the object should remain static is violated in dynamic measurement, making PSP susceptible to object moving, resulting in ripple-like errors in the point clouds. We propose a pixel-wise and frame-wise loopable binomial self-compensation (BSC) algorithm to effectively and flexibly eliminate motion error in the four-step PSP. Our mathematical model demonstrates that by summing successive motion-affected phase frames weighted by binomial coefficients, motion error exponentially diminishes as the binomial order increases, accomplishing automatic error compensation through the motion-affected phase sequence, without the assistance of any intermediate variable. Extensive experiments show that our BSC outperforms the existing methods in reducing motion error, while achieving a depth map frame rate equal to the camera’s acquisition rate (90 fps), enabling high-accuracy 3D reconstruction with a quasi-single-shot frame rate.
@@ 1 Introduction
In recent years, structured light (SL) 3D imaging has been penetrated into diverse fields, e.g., industrial inspection [20] and design [33], medical 3D reconstruction [14, 8, 3], virtual reality [28], and digitalization of cultural heritage [9]. Phase shifting profilometry (PSP) [23], Fourier transform profilometry, binary code [19], and Gray code [22] are famous SL technologies that are widely applied in industry and scientific researches. Among the SL technologies above, PSP is especially favored in high-precision 3D scanning due to its high accuracy, robustness, and pixel-wise property [34]. Many studies have excavated the huge potentiality of PSP in various challenging scenes, such as low signal-to-noise [7, 5], multi-path [31, 30], high reflectivity [18], and global illumination [2, 6, 1], et. al.
@@
In this paper, without depending on any extra intermediate variable, we develop a binomial self-compensation (BSC) method that only utilizes the motion-affected phase sequence itself, thereby effectively and flexibly eliminating the motion error in four-step PSP. To achieve this, first, we design a paraxial binocular structured light system for reconstructing 3D point clouds using a single set of high-frequency fringe patterns. This system intentionally sets a very short baseline between two cameras to achieve a narrow disparity range, facilitating the use of high frequency fringes, while ensuring the baseline between the main camera and projector is long enough for adequate accuracy. Subsequently, we cyclically [32] project high-frequency patterns with $\pi/2$ phase shifting, and directly compute a sequence of phase frames from each four successive captured images. Finally, by adding up a sequence of successive phase frames weighted by binomial coefficients, we achieve self-compensation of motion errors. Our mathematical model demonstrates that motion error exponentially diminishes as the binomial order $K$ increases.
@@ 3 System Setup and Frequency Selection
Figure 1: Diagram of our system.
Our system comprises a main camera, an auxiliary camera, and a projector, as depicted in Fig. 1. Together, the main and auxiliary cameras form a paraxial setup to provide reference depth for phase unwrapping. Simultaneously, the combination of the main camera and projector facilitates precise 3D reconstruction. Following system calibration, we perform stereo rectification on the captured images to align the epipolar lines in the auxiliary system parallelly.
@@ 4 Binomial Self-compensation for Motion Error
where $A^{c}$ is the background intensity, $B^{c}$ is the modulation, $x_{i}$ represents the unknown phase offset induced by the motion from the datum frame to $i$-th frame, note that the datum frame can be arbitrarily chosen, and the true phase value at $i$ th frame is ${\phi_{i}}={\phi_{0}}-i\frac{\pi}{2}$. We choose a $\pi/2$ phase shifting because it only transforms $\sin(\cdot)$ and $\cos(\cdot)$ terms into each other without inducing extra phase offset.
@@
where we define ${\Delta^{\left(0\right)}}{x_{i}}=x_{i}$.
@@
where $\tilde{\phi}_{i}\left(0\right)$ is computed by Eq. (6). We substitute Eq. (8) into Eq. (10) to have
$\begin{split}\tilde{\phi}_{i}^{3}\left(K\right)&=\bar{\phi}_{i}^{3}\left(K\right)-{2^{-K}}{\Delta^{\left({K+2}\right)}}{x_{i+2}}\cos\left({2{\phi_{i}}}\right)\\ &+{2^{-K}}\left({{\Delta^{\left({K+1}\right)}}{x_{i+2}}+{\Delta^{\left({K+1}\right)}}{x_{i+1}}}\right)\sin\left({2{\phi_{i}}}\right)\end{split},$ (11)
@@
$\bar{\phi}_{i}^{3}\left(K\right)={2^{-K}}\sum\limits_{k=0}^{K}{\left[{\left(\begin{split}K\\ k\end{split}\right)\left({{\phi_{i+K-k}}+\delta_{i+K-k}^{3}}\right)}\right]}.$ (12)
@@
$\tilde{\phi}_{i}^{4}\left({K}\right)=\bar{\phi}_{i}^{4}\left({K}\right)+{2^{-K}}\left({{\Delta^{\left({K+1}\right)}}{x_{i}}+{\Delta^{\left({K+1}\right)}}{x_{i+2}}}\right)\cos\left({2{\phi_{i}}}\right),$ (13)
@@
$\bar{\phi}_{i}^{4}\left({K}\right)={2^{-K}}\sum\limits_{k=0}^{K}{\left[{\left(\begin{array}[]{c}K\\ k\end{array}\right)\left({{\phi_{i+K-k}}+\delta_{i+K-k}^{4}}\right)}\right]}.$ (14)
@@
Based on Eq. (11) and (13), it is clear that increasing $K$ produces two main effects: first, the factor $2^{-K}$ exponentially decreases the harmonic amplitudes; second, higher-order difference terms (approach zero as $K$ increases) occur in the harmonic amplitudes. Both effects significantly diminish the motion error.
We measure a flat plate moving towards the camera using our binomial self-compensation (BSC) for motion error, the results visualized in Fig. 3(a) show that the ripples induced by motion in the self-compensated phase $\tilde{\phi}_{i}^{4}\left(K\right)$ reduces significantly with the increment of $K$. $K=0$ represents the raw phase that is affected by motion error. The residual error is caused by the nonlinearity of the binary defocus technique as shown in Fig. 3(b).
We prefer to use four-step phase shifting due to its rapid speed of error convergence as shown in Fig. 3(a), thus, at least five frames of images are needed in BSC, and using eight images achieves a trade-off between accuracy and efficiency in practice. Note that the binomial order $K$ can be flexibly chosen according to the accuracy and time efficiency requirements.
@@
${\phi_{a}}\oplus{\phi_{b}}:=\frac{{{\phi_{a}}+{\phi_{b}}+\pi\left[{{\rm{1+sign}}\left({\left|{{\phi_{a}}-{\phi_{b}}}\right|-\pi}\right)}\right]}}{2},$ (15)
where $\rm{sign}(\cdot)$ denotes the symbolic function. This operation automatically compensates for the phase order error as long as the phase difference is greater than $\pi$. The motion-error-free phase $\tilde{\phi}_{i}\left(K\right)$ can be obtained by employing our binomial self-compensation (BSC), which is summarized as Algorithm 1, where "$\bmod$" represents modular operation. Note that we have corrected the constant phase shifting before adding up two phase frames $\tilde{\phi}_{i}(0)$ and $\tilde{\phi}_{i+1}(0)$. The MATLAB code for implementing the algorithm are given in the supplementary materials.
@@ 6             $\tilde{\phi}_{i}(k+1)\leftarrow\tilde{\phi}_{i}(k)\oplus\tilde{\phi}_{i+1}(k)$
5 Stereo Phase Unwrapping
In the previous section, we obtained the motion-error-free high-frequency phase by conducting BSC, it will be unwrapped in this section. Given the pre-set depth range, the disparity range in the auxiliary camera corresponding to each pixel in the main camera $(x^{m},y^{m})$ is established. Subsequently, a phase-guided sum of absolute difference algorithm [29] is employed to retrieve the correspondence between the main and auxiliary cameras. Further, the projector space coordinates $(x^{p},y^{p})$ can be computed by trifocal tensor or directly re-project the binocular 3D coordinate onto the projector space. Finally, we can have the phase order and compute the unwrapped phase.
@@ 6 Experimental Evaluations
In order to comprehensively demonstrate the performance of our BSC, we conducted four groups of experiments on dynamic objects for evaluating: 1) absolute accuracy of 3D reconstruction, 2) temporal resolution of 3D imaging, 3) robustness on depth discontinuous scenes, and 4) generality on different dynamic objects.
@@ 6.3 Temporal Resolution of 3D Imaging
Our BSC employs cyclic $\pi/2$ phase-shifting fringe patterns, any $K+N$ successive captured images can be used for reconstructing one frame of 3D point clouds. Therefore, our resulting frame rate of 3D point clouds is the same as the camera frame rate, thereby achieving a high temporal resolution of the measurement. That is the frame-wise loopable property we mentioned in the introduction. We successively conducted BSC and $\mu$-FTP to measure a wooden butterfly with wings flapping cyclically, and then manually aligned the timeline, the results are shown in Fig. 8. We noticed that our method can achieve 3D reconstruction with a frame rate of 90 fps (the same as the camera acquisition speed), while $\mu$-FTP reconstructing at 30 fps. Thus, our method is capable of distinguishing tinier motions.
@@ 7 Limitations and Future Outlook
Like most motion compensation algorithms, our BSC assumes low-frequency textures, which may result in artifacts in objects with complex high-frequency textures. Additionally, although our method can generate depth maps at the camera acquisition frame rate, it requires multiple images to compute each depth frame. Consequently, we characterize our approach as providing quasi-single-shot 3D reconstruction, rather than a true single-shot frame rate.
In this paper, we discovered that the motion error can be self-compensated by the characteristics of motion-affected phase frames themselves. However, our BSC is only suitable for a phase shifting step of $\pi/2$. In the future, by employing a divide and conquer strategy [10], we hope to extend this idea to a more general phase-shifting step for suppressing the nonlinear error.
Currently, we restrict the binocular baseline by positioning the cameras side by side, constrained by the physical dimensions of the industrial camera. In subsequent iterations, we plan to improve our prototype by integrating a beam splitter to further reduce the binocular baseline. This adjustment will enable the utilization of fringe patterns with higher frequencies, thereby enhancing measurement accuracy.

### v2

@@ Abstract
Phase shifting profilometry (PSP) is favored in high-precision 3D scanning due to its high accuracy, robustness, and pixel-wise property. However, a fundamental assumption of PSP that the object should remain static is violated in dynamic measurement, making PSP susceptible to object moving, resulting in ripple-like errors in the point clouds. We propose a pixel-wise and frame-wise loopable binomial self-compensation (BSC) algorithm to effectively and flexibly eliminate motion error in the four-step PSP. Our mathematical model demonstrates that by summing successive motion-affected phase frames weighted by binomial coefficients, motion error exponentially diminishes as the binomial order increases, accomplishing automatic error compensation through the motion-affected phase sequence, without the assistance of any intermediate variable. Extensive experiments show that our BSC outperforms the existing methods in reducing motion error, while achieving a depth map frame rate equal to the camera’s acquisition rate (90 fps), enabling high-accuracy 3D reconstruction with a quasi-single-shot frame rate. The code is available at https://github.com/GeyouZhang/BSC.
@@ 1 Introduction
In recent years, structured light (SL) 3D imaging has been penetrated into diverse fields, e.g., industrial inspection [20] and design [33], medical 3D reconstruction [14, 8, 3, lv2023modeling], virtual reality [28], and digitalization of cultural heritage [9]. Phase shifting profilometry (PSP) [23], Fourier transform profilometry, binary code [19], and Gray code [22] are famous SL technologies that are widely applied in industry and scientific researches. Among the SL technologies above, PSP is especially favored in high-precision 3D scanning due to its high accuracy, robustness, and pixel-wise property [34]. Many studies have excavated the huge potentiality of PSP in various challenging scenes, such as low signal-to-noise [7, 5], multi-path [31, 30], high reflectivity [18], and global illumination [2, 6, 1], et. al.
@@
In this paper, without depending on any extra intermediate variable, we develop a binomial self-compensation (BSC) method that only utilizes the motion-affected phase sequence itself, thereby effectively and flexibly eliminating the motion error in four-step PSP. To achieve this, first, we design a paraxial binocular structured light system for reconstructing 3D point clouds using a single set of high-frequency fringe patterns. This system intentionally sets a very short baseline between two cameras to achieve a narrow disparity range, facilitating the use of high frequency fringes, while ensuring the baseline between the main camera and projector is long enough for adequate accuracy. Subsequently, we cyclically [32] project high-frequency patterns with $\pi/2$ phase shifting, and directly compute a sequence of phase frames from each three or four successive captured images. Finally, by adding up a sequence of successive phase frames weighted by binomial coefficients, we achieve self-compensation of motion errors. Our mathematical model demonstrates that motion error exponentially diminishes as the binomial order $K$ increases.
@@ 3 System Setup and Frequency Selection
Figure 1: Diagram of our paraxial binocular structured light system.
Our paraxial binocular structured light system comprises a main camera, an auxiliary camera, and a projector, as depicted in Fig. 1. Together, the main and auxiliary cameras form a paraxial setup to provide reference depth for phase unwrapping. Simultaneously, the combination of the main camera and projector facilitates precise 3D reconstruction. Following system calibration, we perform stereo rectification on the captured images to align the epipolar lines in the auxiliary system parallelly.
@@ 4 Binomial Self-compensation for Motion Error
where $A^{c}$ is the background intensity, $B^{c}$ is the modulation, $x_{i}$ represents the unknown phase offset induced by the motion from the datum frame to $i$-th frame, note that the datum frame can be arbitrarily chosen, and the true phase value at $i$ th frame is ${\phi_{i}}={\phi_{0}}-i\frac{\pi}{2}$. We choose a $\pi/2$ phase shifting as it only transforms $\sin(\cdot)$ and $\cos(\cdot)$ terms into each other without inducing extra phase offset.
@@
where we define ${\Delta^{\left(0\right)}}{x_{i}}=x_{i}$. Eq. (5) is the key to motion error compensation, and we will use it later.
@@
where $\tilde{\phi}_{i}\left(0\right)$ is computed by Eq. (6). We substitute Eq. (5) and (8) into Eq. (10) to have
$\begin{split}\tilde{\phi}_{i}^{3}\left(K\right)&=\bar{\phi}_{i}^{3}\left(K\right)-{2^{-\left(K+2\right)}}{\Delta^{\left({K+2}\right)}}{x_{i+2}}\cos\left({2{\phi_{i}}}\right)\\ &+{2^{-\left(K+2\right)}}\left({{\Delta^{\left({K+1}\right)}}{x_{i+2}}+{\Delta^{\left({K+1}\right)}}{x_{i+1}}}\right)\sin\left({2{\phi_{i}}}\right)\end{split},$ (11)
@@
$\bar{\phi}_{i}^{3}\left(K\right)={2^{-\left(K+2\right)}}\sum\limits_{k=0}^{K}{\left[{\left(\begin{split}K\\ k\end{split}\right)\left({{\phi_{i+K-k}}+\delta_{i+K-k}^{3}}\right)}\right]}.$ (12)
@@
$\tilde{\phi}_{i}^{4}\left({K}\right)=\bar{\phi}_{i}^{4}\left({K}\right)+{2^{-\left(K+2\right)}}\left({{\Delta^{\left({K+1}\right)}}{x_{i}}+{\Delta^{\left({K+1}\right)}}{x_{i+2}}}\right)\cos\left({2{\phi_{i}}}\right),$ (13)
@@
$\bar{\phi}_{i}^{4}\left({K}\right)={2^{-\left(K+2\right)}}\sum\limits_{k=0}^{K}{\left[{\left(\begin{array}[]{c}K\\ k\end{array}\right)\left({{\phi_{i+K-k}}+\delta_{i+K-k}^{4}}\right)}\right]}.$ (14)
@@
Based on Eq. (11) and (13), it is clear that increasing $K$ produces two main effects: first, the factor $2^{-(K+2)}$ exponentially decreases the harmonic amplitudes; second, higher-order difference terms (approach zero as $K$ increases) occur in the harmonic amplitudes. Both effects significantly diminish the motion error.
We measure a flat plate moving towards the camera using our binomial self-compensation (BSC) for motion error, the results visualized in Fig. 3(a) show that the ripples induced by motion in the self-compensated phase $\tilde{\phi}_{i}^{4}\left(K\right)$ reduces significantly with the increment of $K$. $K=0$ represents the raw phase that is affected by motion error. The residual error is caused by the nonlinearity of the binary defocus technique as shown in Fig. 3(b). At least four frames of images are needed in BSC, and using eight images achieves a trade-off between accuracy and efficiency in practice. Note that the binomial order $K$ can be flexibly chosen according to the accuracy and time efficiency requirements.
@@
${\phi_{a}}\oplus{\phi_{b}}:=\left\{\begin{array}[]{l}0.5\left({{\phi_{a}}+{\phi_{b}}}\right),\left|{{\phi_{a}}-{\phi_{b}}}\right|>\pi\\ 0.5\left({{\phi_{a}}+{\phi_{b}}}\right)+\pi,\left|{{\phi_{a}}-{\phi_{b}}}\right|<\pi\end{array}\right..$ (15)
This operation automatically compensates for the phase order error as long as the phase difference is greater than $\pi$. The motion-error-free phase $\tilde{\phi}_{i}\left(K\right)$ can be obtained by employing our binomial self-compensation (BSC), which is summarized as Algorithm 1, where "$\bmod$" represents modular operation. Note that we have corrected the constant phase shifting before adding up two phase frames $\tilde{\phi}_{i}(0)$ and $\tilde{\phi}_{i+1}(0)$. The MATLAB code for implementing the algorithm are given in the github repository.
5 Stereo Phase Unwrapping
In the previous section, we obtained the motion-error-free high-frequency phase by conducting BSC. Subsequently, we employ a sum of absolute difference algorithm on the wrapped phase to retrieve the correspondence between the main and auxiliary cameras within the disparity range obtained by pre-set depth range. Further, the projector space coordinates $(x^{p},y^{p})$ can be computed by directly re-projecting the binocular 3D coordinate onto the projector space. Finally, we can have the phase order and compute the unwrapped phase.
@@ 6 Experimental Evaluations
In order to comprehensively demonstrate the superiority of our BSC, we conducted four groups of experiments on dynamic objects for evaluating: 1) absolute accuracy of 3D reconstruction, 2) temporal resolution of 3D imaging, 3) robustness on depth discontinuous scenes, and 4) generality on different dynamic objects.
@@ 6.3 Temporal Resolution of 3D Imaging
Our BSC employs cyclic $\pi/2$ phase-shifting fringe patterns, any $K+N$ successive captured images can be used for reconstructing one frame of 3D point clouds. Therefore, our resulting frame rate of 3D point clouds is the same as the camera frame rate, thereby achieving a high temporal resolution of the measurement. That is the frame-wise loopable property we mentioned in the introduction. We successively conducted BSC and $\mu$-FTP to measure a wooden butterfly with wings flapping cyclically, and then manually aligned the timeline, the results are shown in Fig. 8. We noticed that our method can achieve 3D reconstruction with a frame rate of 90 fps (the same as the camera frame rate), while $\mu$-FTP reconstructing at 30 fps. Thus, our method is capable of distinguishing tinier motions.
@@ 7 Limitations and Future Outlook
Figure 10: 3D error VS inter-frame phase-shifting error $|\Delta x_{i}|$.
@@
High-frequency textures: our BSC assumes low-frequency textures like most motion compensation algorithms, it is robust against moderate contrast textures. However, when dealing with objects that have high contrast textures, our method induces artifacts similar to those seen with typical phase-shifting methods.
Phase unwrapping error: we adopt a common block matching algorithm for stereo correspondence, inducing obivious phase unwrapping error when measuring complex surfaces. In the future, we will conduct a hierarchical adjustment for phase outliers to optimize the accuracy of the point cloud.
Applicable speed range: Our BSC is predicated on the assumption that the phase-shift error $x_{i}$ is small. Consequently, as shown in Fig. 10, the residual motion error arises as $|\Delta x_{i}|$ increases. Properly reducing the fringe frequency could decrease $|\Delta x_{i}|$, alleviating the performance loss of BSC for high-speed objects.